# On the Limits of Self-Improving in LLMs and Why AGI, ASI and the Singularity Are Not Near Without Symbolic Model Synthesis

### v1

@@ Abstract
We formalise recursive self-training in Large Language Models (LLMs) and Generative AI as a discrete-time dynamical system and prove that, as training data become increasingly self-generated ($\alpha_{t}\to 0$), the system undergoes inevitably degenerative dynamics. We derive two fundamental failure modes: (1) Entropy Decay, where finite sampling effects cause a monotonic loss of distributional diversity (mode collapse), and (2) Variance Amplification, where the loss of external grounding causes the model’s representation of truth to drift as a random walk, bounded only by the support diameter. We show these behaviours are not contingent on architecture but are consequences of distributional learning on finite samples. We further argue that Reinforcement Learning with imperfect verifiers suffers similar semantic collapse. To overcome these limits, we propose a path involving symbolic regression and program synthesis guided by Algorithmic Probability. The Coding Theorem Method (CTM) allows for identifying generative mechanisms rather than mere correlations, escaping the data-processing inequality that binds standard statistical learning. We conclude that while purely distributional learning leads to model collapse, hybrid neurosymbolic approaches offer a coherent framework for sustained self-improvement.
@@ 1 Introduction
The notion of a technological or AI Singularity, popularised by Vernor Vinge and Ray Kurzweil (12), posits a future inflection point where artificial intelligence surpasses human intellect, leading to an “intelligence explosion” of unforeseeable consequence (23; 13). Central to this hypothesis is the concept of recursive self-improvement: an AI system with the capacity to inspect and enhance its own architecture or training processes would initiate a positive feedback loop, with each generation of the AI being more intelligent than the last, leading to exponential growth in its capabilities.
@@ 2.2 Mutual information, DPI and contraction inequalities
2.3 Algorithmic complexity, information and probability
@@ 2.4 The Coding Theorem and Block Decomposition Methods (CTM and BDM)
2.5 Algorithmic Information Dynamics (AID)
@@ 2.6 Neurosymbolic operators
We define a one-step update as a composition of operators:
@@ 3.3 Model Collapse: The Curse of Degenerative Recursion
Our work aims to formalise these observations into a rigorous proof of convergence, generalising the argument to any generative model and directly linking this inevitable collapse to the impossibility of a GenAI-driven Singularity-type argument based on current mainstream architectures.
@@ 4 Self-Referential Training as a Dynamical System
The transition from model $M_{t}$ to $M_{t+1}$ occurs via a training process on a new dataset. This dataset is a mixture of authentic and synthetic data.
  • Let $\alpha_{t}\in[0,1]$ be the proportion of new, authentic data drawn from $P$ at iteration $t$.
  • The remaining proportion, $1-\alpha_{t}$, consists of synthetic data sampled from the current model’s distribution, $Q_{t}$.
  • The training distribution for the next model, $P_{t}^{\prime}$, is a convex combination:
    $P_{t}^{\prime}=\alpha_{t}P+(1-\alpha_{t})Q_{t}$ (9)
  • The new model $M_{t+1}$ is obtained by optimising its parameters to minimise the divergence between its distribution $Q_{t+1}$ and the training distribution $P_{t}^{\prime}$. A common objective is to minimise the KL-divergence:
    $Q_{t+1}=\arg\min_{Q\in\mathcal{Q}}D_{KL}(P_{t}^{\prime}\,||\,Q)$ (10)
    This is equivalent to maximising the log-likelihood of the data sampled from $P_{t}^{\prime}$.
@@ 4.4 The Realistic Case ($\alpha_{t}\to 0$)
The parameter $\alpha_{t}\in[0,1]$ denotes the proportion of fresh, externally grounded data drawn from the true distribution $P$ at iteration $t$. The condition $\alpha_{t}\to 0$ means $\lim_{t\to\infty}\alpha_{t}=0$, i.e. for every $\varepsilon>0$ there exists $T$ such that $\alpha_{t}<\varepsilon$ for all $t\geq T$. Equivalently, the training mixture
@@
becomes asymptotically self-referential, with the influence of $P$ vanishing in the limit. This regime formalises the autonomy requirement implicit in Singularity-style recursive self-improvement arguments: the system must eventually rely predominantly on its own outputs rather than on a persistent external oracle.
@@ 5 Implications for AGI, ASI and Singularity arguments
One core premise of arguments around Artificial General Intelligence (AGI), Artificial Super Intelligence (ASI) and the so-called AI ‘Singularity’ is a recursive process whereby unbounded growth is possible without external or human intervention. Our analysis shows that the self-referential training process, far from being unbounded, is a convergent process. It leads not to an “explosion” but to an “implosion” of informational diversity and knolwedge inference.
  • Convergence vs. Divergence: A Singularity requires a divergent process, where capability grows exponentially. Model collapse is a convergent process, where the model’s distribution approaches a static, limited fixed point.
  • The Necessity of Grounding: The theorems demonstrate that without a persistent connection to an external, truthful data source ($P$), an AI system’s model of reality will inevitably drift and degrade. The value $\alpha>0$ is not a temporary bootstrap but a permanent necessity. AGI (9) cannot be developed in a closed box; it requires continuous, active grounding in the real world.
  • Errors are Amplified, Not Corrected: In a self-referential loop, any inaccuracies, biases, or “hallucinations” present in model $Q_{t}$ become codified as “truth” in the training data for model $Q_{t+1}$. The system lacks an external error correction mechanism and will progressively amplify its own flaws.
5.1 A Fundamental Kantian Distinction: Analytic vs. Synthetic AI
To understand the philosophical limits of GenAI, we can borrow from Immanuel Kant’s epistemology. Kant distinguished between two types of judgements:
  • Analytic Judgements: The predicate is contained within the concept of the subject (e.g., “All bachelors are unmarried”). These judgements are explicative; they do not add new knowledge but clarify what is already known via logical implication.
  • Synthetic Judgements: The predicate is not contained within the concept of the subject (e.g., “The cat is on the mat” or “7+5=127+5=12”). These require empirical observation or, crucially, active computation to verify.
We can apply this framework to AI. Current Generative AI is a fundamentally analytic system. It is trained on a massive, but finite, dataset representing a snapshot of human knowledge ($P$). Its operations consist of identifying patterns, correlations, and structures within this dataset and then interpolating or recombining them to generate outputs. While the outputs may be surprising, they are ultimately derivative of the correlations latent within the training data (the “subject”).
A true AGI, and certainly one capable of initiating a Singularity, must be capable of making synthetic judgements. In a computational context, this corresponds to execution rather than prediction. An LLM predicts the output of a Python script based on its training data (analytic); a Symbolic engine runs the script (synthetic). The latter generates a result that is not statistically entailed by the input but is the result of an irreducible computational process.
Model collapse is the mathematical manifestation of this philosophical limit. When an analytic engine is forced to feed on its own outputs, it has nothing new to analyse. It can only re-process and re-combine its existing knowledge, leading to a caricature of its former self. It cannot synthesise new knowledge because it lacks the mechanism of computational verification or external grounding, and therefore it cannot “improve” in any meaningful sense.
@@ 5.2 Causality and Neurosymbolic Approaches as a Path Beyond Collapse
Thus, KL divergence provides a unifying mathematical framework that encompasses the most common objectives in deep learning. Our collapse proofs therefore apply broadly to existing transformer and generative model training regimes.
@@ 5.3 Statistical Loss Functions Contribute to Collapse
where $M(\cdot)$ denotes a causal representation of a distribution, for example, a structural causal model or a symbolic regression (24; 26), then the system optimises not for correlation but for invariance under intervention. In this framework, even synthetic data can yield new information if they encode counterfactual or interventional predictions. Unlike KL-based training, which contracts toward fixed points, causal objectives permit the creation of genuinely synthetic knowledge, in line with Kant’s distinction between analytic and synthetic judgments.
@@ 5.6.3 Refining the Contraction
The algorithmic update ensures that $I(M;Q_{t+1}^{alg})\geq I(M;Q_{t}^{stat})$, explicitly escaping the data-processing inequality (as per Proposition 9) by injecting the universal prior $m$. This allows the system to maintain alignment with the ground truth mechanism $M$ even when the external data fraction $\alpha_{t}\to 0$, provided the underlying reality $M$ remains algorithmically simple.
@@ 6 Conclusion
We have demonstrated that the “bootstrap” hypothesis of AGI—that an AI system can recursively improve itself indefinitely using its own outputs—is mathematically unsound under current distributional learning paradigms. Specifically, we proved that the self-referential loop constitutes a supermartingale with respect to entropy, leading inevitably to information loss and mode collapse. Furthermore, in the absence of a persistent external grounding signal ($\alpha>0$), the model’s centroid drifts from the true distribution via a random walk driven by optimisation noise.
Critically, this limitation extends to Reinforcement Learning systems relying on learned or proxy verifiers, which are themselves subject to semantic collapse. The system is information-theoretically closed: it can re-represent existing errors but cannot generate new information about the environment.
The only viable path to escape these attractor dynamics is to move from analytic prediction (correlation-based) to synthetic computation (mechanism-based). By integrating statistical learning with algorithmic information theory—specifically the Coding Theorem Method (CTM) and Block Decomposition Method (BDM)—we can construct objectives that penalise causal incoherence rather than just distributional divergence. Future work must focus on these neurosymbolic architectures, where “improvement” is defined not by lower perplexity, but by the discovery of more concise, algorithmically probable programs that explain the observed data.

### v2

# On the Limits of Self-Improving in Large Language Models: The Singularity Is Not Near Without Symbolic Model Synthesis

@@ Abstract
We formalise recursive self-training in Large Language Models (LLMs) and Generative AI as a discrete-time dynamical system. We prove that if the proportion of exogenous, externally grounded signal $\alpha_{t}$ vanishes asymptotically ($\alpha_{t}\to 0$), the system undergoes degenerative dynamics. We derive two fundamental failure modes: (1) Entropy Decay, where finite sampling effects induce monotonic loss of distributional diversity, and (2) Variance Amplification, where the absence of persistent grounding causes distributional drift via a random-walk mechanism. These behaviours are architectural invariants of distributional learning on finite samples. We show that the collapse results apply specifically to closed-loop density matching without persistent external signal. Systems with non-vanishing exogenous grounding fall outside this regime. However, mainstream Singularity, AGI, and ASI narratives typically posit systems that become increasingly autonomous and require little to no human or external intervention for self-improvement. In that autonomy regime, the vanishing-signal condition is satisfied, and collapse follows under KL-based objectives. To overcome these limits, we propose neurosymbolic integration based on algorithmic probability and program synthesis. The Coding Theorem Method (CTM) enables identification of generative mechanisms rather than mere correlations, escaping the distribution-only constraints that bind standard statistical learning. We conclude that fully autonomous recursive density matching leads to degenerative fixed points, whereas externally anchored or mechanism-based approaches operate under fundamentally different asymptotic dynamics.
@@ 1 Introduction
The notion of a technological or AI Singularity, popularised by Vernor Vinge and Ray Kurzweil (14), posits a future inflection point where artificial intelligence surpasses human intellect, leading to an “intelligence explosion” of unforeseeable consequence (25; 15). Central to this hypothesis is the concept of recursive self-improvement: an AI system with the capacity to inspect and enhance its own architecture or training processes would initiate a positive feedback loop, with each generation of the AI being more intelligent than the last, leading to exponential growth in its capabilities.
@@
Recent empirical work on model collapse has clarified that degenerative dynamics depend critically on the proportion of synthetic to real data. While early discussions considered scenarios in which synthetic data progressively replace real data, more recent studies show that collapse can be mitigated when real data continue to accumulate at a sufficient rate (2). In the formalism developed in this paper, this corresponds to maintaining $\inf_{t}\alpha_{t}>0$, where $\alpha_{t}$ denotes the proportion of exogenous, externally grounded signal.
The central question for Singularity, AGI, and ASI arguments is therefore not whether collapse can be avoided under persistent external grounding, but whether recursive self-improvement is assumed to become increasingly autonomous. Strong versions of these hypotheses typically posit systems that require progressively less human or environmental input in order to improve themselves. Such autonomy implies that the relative contribution of externally grounded signal becomes negligible over time, i.e. that $\alpha_{t}\to 0$. It is precisely this autonomy regime that our theorems analyse.
@@ 2 Preliminaries and Notation
2.1 Coding and Block Decomposition Methods and Algorithmic Information Dynamics
Kolmogorov complexity and algorithmic probability provide a formal framework for measuring information in terms of generative mechanisms rather than statistical frequencies. Given a fixed universal Turing machine, the Kolmogorov complexity of a finite object is defined as the length of the shortest programme that produces that object as output. It therefore quantifies information as minimal description length in programme space. An object is considered algorithmically simple if it admits a short generating programme, and algorithmically random if no description significantly shorter than the object itself exists.
Algorithmic probability complements this notion by assigning to each object the probability that a randomly generated programme will produce it. Short programmes contribute more weight than long ones, yielding a universal prior that favours simpler generative explanations. The coding theorem links the two concepts: objects with high algorithmic probability have low Kolmogorov complexity, and vice versa. Together, these notions shift the focus from observed statistical regularities to the underlying mechanisms capable of generating the data, forming the basis for a mechanistic interpretation of information beyond Shannon entropy.
@@ 2.1 Coding and Block Decomposition Methods and Algorithmic Information Dynamics
Algorithmic Information Dynamics (AID) is a theoretical framework that studies the structure, evolution and causal content of systems through the lens of Algorithmic Information Theory. Rather than relying on statistical regularities, it characterises patterns, randomness and complexity in terms of the length of the shortest effective description capable of generating an object or process.
The central idea is that the behaviour of a system can be understood by analysing how its algorithmic information content changes under perturbations. By estimating contributions of individual elements to the overall description length, the framework aims to distinguish structural, causal components from incidental or noisy ones. In this way, Algorithmic Information Dynamics provides tools for inferring generative mechanisms, identifying causal pathways and quantifying the directionality of information flow in deterministic and stochastic systems alike.
Although rooted in uncomputable quantities such as Kolmogorov complexity, the framework employs computable approximations and decomposition methods to make empirical analysis possible across domains including dynamical systems, networks and biological processes.
@@ 2.2 Neurosymbolic operators
A neurosymbolic operator is a transformation that combines two distinct modes of inference:
  • Subsymbolic / statistical learning, typically implemented by neural networks that approximate probability distributions from data.
  • Symbolic / structural reasoning, which operates over discrete representations such as rules, equations, graphs, logical constraints, or programmes.
A neurosymbolic operator refers to a transformation that integrates statistical learning with structured, rule-based or mechanistic reasoning within a single update process. Modern neural systems are highly effective at modelling complex probability distributions from large datasets, capturing high-dimensional correlations and producing flexible interpolations. However, such systems operate primarily in distribution space: they adjust parameters to minimise a loss function defined over observed frequencies, without explicit access to underlying structural constraints or generative mechanisms.
Symbolic systems, by contrast, manipulate discrete representations such as logical rules, equations, causal graphs, or programmes. They operate in structure space, enforcing invariants, compositional rules, and mechanistic relations that are not reducible to correlation alone. While neural models excel at approximation, symbolic systems excel at constraint, explanation, and generalisation under transformation.
A neurosymbolic operator combines these two regimes. Formally, it updates a learned model not only to improve statistical fit to data, but also to satisfy or approximate a set of structural constraints derived from symbolic representations. The operator therefore acts simultaneously in distribution space and in structure space. In practice, this may take the form of projecting a neural model onto a constraint set defined by logical rules, causal relations, conservation laws, or syntactic grammars, or of guiding model updates through programme or equation synthesis consistent with the data.
In an algorithmic framework, the notion extends further: the symbolic component may consist of candidate generative programmes weighted by algorithmic probability, and updates are evaluated not solely by reduction in statistical divergence but by their coherence under structured perturbations. In this setting, a neurosymbolic operator is not merely a hybrid architecture but a principled learning mechanism that integrates statistical evidence with mechanistic structure, thereby moving beyond purely correlational inference.
The intuition is that neural models are powerful at interpolation and pattern extraction in high-dimensional spaces, while symbolic systems are powerful at enforcing invariants, causal structure, compositionality, and mechanistic constraints. A neurosymbolic operator acts as a bridge: it updates or transforms a model by jointly optimising statistical fit and structural consistency.
We can define a one-step update as a composition of operators:
@@ 3.2 Model Collapse: The Curse of Degenerative Recursion
Our work formalises these observations into a rigorous proof of convergence under the vanishing-exogenous-signal condition ($\alpha_{t}\to 0$), generalising the argument to any generative model trained via distributional objectives. Empirical studies indicate that collapse depends critically on the proportion of real versus synthetic data, with degradation emerging when synthetic data dominate and stability maintained when sufficient external signal persists (2). Our analysis provides the formal condition under which such degradation becomes inevitable.
@@ 4 Self-Referential Training as a Dynamical System
The transition from model $M_{t}$ to $M_{t+1}$ occurs via training on a mixture of exogenous and endogenous signal.
  • Let $\alpha_{t}\in[0,1]$ be the proportion of exogenous training signal at iteration $t$, meaning signal whose correctness or generating process is determined independently of the model’s current distribution. This may include fresh samples drawn from $P$, environmental feedback, simulator outcomes, or outputs validated by an external verifier.
  • The remaining proportion, $1-\alpha_{t}$, consists of endogenous samples generated from the current model distribution $Q_{t}$ (or an ensemble derived from it).
  • The training distribution for the next model, $P_{t}^{\prime}$, is the convex combination:
    $P_{t}^{\prime}=\alpha_{t}P+(1-\alpha_{t})Q_{t}.$ (9)
  • The new model $M_{t+1}$ is obtained by optimising its parameters to minimise divergence between its distribution $Q_{t+1}$ and the training distribution $P_{t}^{\prime}$. A common objective is to minimise KL divergence:
    $Q_{t+1}=\arg\min_{Q\in\mathcal{Q}}D_{KL}(P_{t}^{\prime}\,||\,Q).$ (10)
    This is equivalent to maximising the log-likelihood of data sampled from $P_{t}^{\prime}$.
@@ 4.4 The Realistic Case ($\alpha_{t}\to 0$)
Theorem 5 (Scope Condition: Collapse Requires Vanishing Exogenous Signal).
Let $\{Q_{t}\}$ evolve under the update rule
$Q_{t+1}=\arg\min_{Q\in\mathcal{Q}}D_{KL}(P_{t}^{\prime}\,||\,Q),\quad P_{t}^{\prime}=\alpha_{t}P+(1-\alpha_{t})Q_{t}.$
If $\inf_{t}\alpha_{t}>0$ (i.e. the exogenous signal remains persistently bounded away from zero), and the model family $\mathcal{Q}$ has sufficient capacity, then in the idealised infinite-capacity limit the sequence $\{Q_{t}\}$ converges to $P$.
Conversely, the degenerative dynamics proven in Theorems 2 and the variance amplification result require $\alpha_{t}\to 0$.
Proof.
If there exists $\alpha_{*}>0$ such that $\alpha_{t}\geq\alpha_{*}$ for all $t$, then the update operator contains a persistent contraction term toward $P$. Under the infinite-capacity idealisation, the recursion reduces to
$Q_{t+1}=\alpha_{t}P+(1-\alpha_{t})Q_{t},$
which forms a weighted average with a non-vanishing restoring force. Standard contraction arguments imply convergence to $P$. The collapse results require that this restoring force vanish asymptotically, i.e. $\alpha_{t}\to 0$, in which case the system becomes asymptotically self-referential. ∎
The parameter $\alpha_{t}\in[0,1]$ denotes the proportion of exogenous training signal available at iteration $t$, that is, signal whose truth value or generating process is determined independently of the model’s current distribution. In the simplest instantiation used in this paper, this exogenous component is represented by fresh samples from the true distribution $P$.
The condition $\alpha_{t}\to 0$ means $\lim_{t\to\infty}\alpha_{t}=0$, i.e. for every $\varepsilon>0$ there exists $T$ such that $\alpha_{t}<\varepsilon$ for all $t\geq T$. Equivalently, the training mixture
@@
becomes asymptotically self-referential, with the influence of exogenous signal vanishing in the limit.
This regime formalises the autonomy requirement implicit in certain Singularity-style recursive self-improvement arguments: the system must eventually rely predominantly on its own outputs rather than on a persistent external evaluator, simulator, or data-generating process. The scope clarifications in Section 4.9 make explicit that externally anchored self-improvement (fixed axioms and invariant evaluators) does not satisfy this vanishing-exogenous-signal assumption, even when data are produced via self-play or simulation.
@@ 4.5 Extension to Multi-Modal Ensembles
4.6 Closed-Loop Training vs Externally Anchored Optimisation
The model collapse results presented above apply to a specific structural regime. It is therefore essential to distinguish between two fundamentally different forms of recursive training.
1. Closed-Loop Density Matching.
In the collapse setting, the model is iteratively trained to approximate a distribution that increasingly consists of its own samples. When $\alpha_{t}\to 0$, the update becomes:
$Q_{t+1}\approx\arg\min_{Q}\mathbb{E}_{x\sim Q_{t}}[-\log Q(x)].$
This defines a self-referential transformation on the space of distributions.
In this regime:
  • The objective is implicit in the data distribution itself.
  • There is no external correctness criterion beyond likelihood.
  • The optimisation landscape is shaped by $Q_{t}$ rather than by an independently specified target.
The collapse theorem characterises precisely this regime.
2. Externally Anchored Optimisation.
By contrast, many canonical examples of “self-improvement” operate under additional structural assumptions:
  1. 1.
    A fixed axiomatic base (e.g., game rules or physical laws).
  2. 2.
    An externally defined objective or verifier.
  3. 3.
    A bounded task domain in which improvement is measured.
In such systems, even if training data are generated through self-play or simulation, the evaluation signal is not defined by the model’s own distribution.
Instead, the update operator takes the form:
$Q_{t+1}=\mathcal{T}(Q_{t};\mathcal{E}),$
where $\mathcal{E}$ denotes an external evaluator or environment.
The asymptotic behaviour of this operator is governed by properties of the external objective, not by self-consistency alone.
The collapse results apply to closed-loop generative retraining without persistent exogenous signal. They do not automatically extend to systems in which optimisation remains externally grounded.
This distinction is crucial when drawing conclusions about AGI or recursive self-improvement.
@@ 5 Implications for AGI, ASI and Singularity Arguments
The results established in this paper apply to a specific structural regime: recursive generative retraining in which the proportion of exogenous signal $\alpha_{t}$ vanishes asymptotically.
Under this regime, the system becomes information-theoretically closed. Learning reduces to self-consistency under a distributional objective, and the update operator converges to a fixed point determined by its own prior state. Theorems 2 and the variance amplification result show that in this regime entropy decays and distributional drift accumulates.
The theorems establish that:
  • Purely self-referential density matching ($\alpha_{t}\to 0$) induces contraction toward a degenerate or distorted stationary distribution.
  • Mutual information with the true generating distribution $P$ cannot increase under distribution-only updates.
  • Autonomy in the strong sense of vanishing external grounding leads to asymptotic self-reference rather than unbounded improvement.
The results do not prove that all forms of recursive self-improvement collapse.
If $\inf_{t}\alpha_{t}>0$, meaning the system receives persistent exogenous signal, then the contraction toward $P$ remains active. Systems operating under fixed axioms, externally defined objectives, or invariant verifiers (e.g. formally specified environments) do not satisfy the $\alpha_{t}\to 0$ condition.
Such systems may converge to task-optimal stationary distributions rather than degenerate measures. This distinction is consistent with classical results in stochastic approximation theory, where convergence under persistent excitation is guaranteed when the update receives non-vanishing external signal (1). The collapse theorems proven here apply specifically to the complementary regime in which the excitation term vanishes asymptotically ($\alpha_{t}\to 0$).
Therefore, any claim that a distributional generative model will achieve unbounded recursive improvement must specify whether its learning signal becomes endogenised. If the system asymptotically relies on its own outputs without persistent grounding, collapse follows under the assumptions of this paper. If improvement instead remains externally anchored, the asymptotic regime is fundamentally different and must be analysed under persistent-excitation dynamics.
The Singularity hypothesis, when framed as fully autonomous recursive density matching, is mathematically incompatible with sustained growth under current KL-based generative learning paradigms.
@@ 5.1 Causality and Neurosymbolic Approaches as a Path Beyond Collapse
Thus, KL divergence provides a unifying mathematical framework that encompasses the most common objectives in deep learning. Our collapse proofs, therefore, apply broadly to existing transformer and generative model training regimes.
@@ 5.2 Statistical Loss Functions Contribute to Collapse
where $M(\cdot)$ denotes a causal representation of a distribution, for example, a structural causal model or a symbolic regression (26; 28), then the system optimises not for correlation but for invariance under intervention. In this framework, even synthetic data can yield new information if they encode counterfactual or interventional predictions. Unlike KL-based training, which contracts toward fixed points, causal objectives permit the creation of genuinely synthetic knowledge, in line with Kant’s distinction between analytic and synthetic judgments.
@@ 5.5.3 Refining the Contraction
The algorithmic update ensures that $I(M;Q_{t+1}^{alg})\geq I(M;Q_{t}^{stat})$, explicitly escaping the data-processing inequality (as per Proposition 10) by injecting the universal prior $m$. This allows the system to maintain alignment with the ground truth mechanism $M$ even when the external data fraction $\alpha_{t}\to 0$, provided that the underlying reality $M$ remains algorithmically simple.
@@ 6 Conclusion
We have formalised recursive self-training in generative models as a discrete-time dynamical system and derived two structural failure modes that arise when the proportion of exogenous signal $\alpha_{t}$ vanishes asymptotically. Under this autonomy condition, the update operator becomes increasingly self-referential. Entropy decays due to finite-sample contraction, and distributional drift accumulates through variance amplification. The system converges not toward expanding intelligence, but toward stationary fixed points determined by its own prior state.
The impossibility result is conditional rather than universal. It applies to recursive density matching under KL-type objectives when the influence of externally grounded signal disappears in the limit. If $\alpha_{t}\to 0$, the system becomes information-theoretically closed: no new information about the generating distribution can be created through self-sampling alone. Mutual information with the true distribution cannot increase, and optimisation reduces to self-consistency rather than improvement relative to the world.
This conditional structure is crucial for evaluating claims about AGI, ASI, and the Singularity. Strong versions of these hypotheses typically posit systems that become increasingly autonomous and require little to no human or environmental intervention in order to improve themselves. Such autonomy implicitly entails a regime in which exogenous signal becomes negligible relative to internally generated data. In that regime, our theorems apply directly. Recursive self-improvement framed as progressively self-contained generative retraining cannot yield unbounded growth under standard distributional learning dynamics.
By contrast, systems that retain persistent external grounding ($\inf_{t}\alpha_{t}>0$), operate under fixed axioms with invariant evaluators, or optimise against formally specified environments fall outside the collapse regime proven here. Their asymptotic behaviour must be analysed under persistent-excitation dynamics rather than closed-loop contraction. Our results therefore do not rule out improvement in externally anchored systems; they rule out fully autonomous recursive density matching as a path to indefinite intelligence growth.
The deeper limitation is structural. KL-based and related divergence objectives optimise correlations within observed distributions but do not recover generative mechanisms. When such objectives dominate the update rule, recursive retraining contracts distributional diversity and amplifies internal biases. To escape these attractor dynamics, learning must move beyond analytic correlation toward synthetic mechanism discovery.
Neurosymbolic integration grounded in algorithmic information theory provides a principled direction. By incorporating program synthesis, algorithmic probability, and causal perturbation analysis through tools such as CTM and BDM, the learning objective can shift from distributional fit to mechanistic coherence. In such a framework, improvement is measured not by lower perplexity alone but by convergence toward simpler, causally invariant generative programmes consistent with the data.
In summary, fully autonomous recursive generative retraining under standard statistical objectives leads to degenerative fixed points rather than intelligence explosion. Sustained self-improvement requires persistent grounding or a transition from distributional optimisation to mechanism-based inference. The Singularity, if it is to be coherent, cannot be built upon closed-loop density matching alone.